OpenStudy (texmechanica):

What is the slope of the line represented by the equation y = x/4 - 6? A. -6 B. 4 C. 3/2 D. 1/4

OpenStudy (texmechanica):

@Vijeya3 Please help, I stink at slopes.

OpenStudy (vijeya3):

lol...do u knw the meaning of this equation?

OpenStudy (texmechanica):

No.

OpenStudy (vijeya3):

Okay...there is a formula... y=fx+c where c is the intercept f is the slope and x is the x-coordinate and y is the y-coordinate Does this give u any idea how to solve it?

OpenStudy (texmechanica):

I think so yeah.

OpenStudy (vijeya3):

So..what do u think is the answer? To be honest...I have learnt this thing this very morning!:D

OpenStudy (anonymous):

I thought slope was "y=mx+b"

OpenStudy (vijeya3):

@Maria195 what's in an alphabet...My textbook says it is f...ur textbook says that it is m...:D

OpenStudy (anonymous):

Lol I thought that's how it always was. You can even look up "slope-intercept form" on Google, it'll come up with the same thing.

OpenStudy (vijeya3):

Okay...i dunno...my textbook says whatever i have written...but the idea is the same right?

563blackghost (563blackghost):

You use the formula.... \(\huge\color{green}{y=mx+b}\) \(\huge{~~~~~~~~|~~~~~~~~~|}\) \(\Large\color{red}{~~~~~~~~slope~~~~~y-intercept}\)

563blackghost (563blackghost):

They are the same @Vijeya3 :P

OpenStudy (vijeya3):

Sure....ofcourse it is..:D

OpenStudy (anonymous):

But I think the answer might be D

OpenStudy (texmechanica):

I think so as well.

563blackghost (563blackghost):

That is correct ^^ It is infact \(\huge\color{red}{\frac{1}{4}}\)

OpenStudy (texmechanica):

Thanks guys.

563blackghost (563blackghost):

\(\Huge\cal\rlap{\color{red}{No~Problem!:)}}{\color{blue}{\;No~Problem!:)}}\)